# BlackboxBench: A Comprehensive Benchmark of Black-box Adversarial Attacks

### v1

@@ IV-D2 Input dimensionality
We now quantify the sensitivity to input dimensionality via conducting untargeted, $\ell_{\infty}$ attack on ImageNet[41] dataset with varying image size. It’s important to note that when working with a clean dataset with downsampled images, the clean accuracy naturally decreases, potentially confusing our analysis. Therefore, we use the improvement in ASR after adding adversarial perturbations on benign examples resized to different dimensionalities. This metric only serves as the measure of the transfer-based attack performance. Query-based attacks define the percentage of finally adversarial images in initially non-adversarial images as ASR and thus are not influenced by the dropped clean accuracy. Fig. 9 shows that all adversaries suffer from a considerable failure rate when facing the robustness existing in low dimensional classification tasks. This intuition behind this phenomenon is clearly illustrated in Fig. 10(a). Consider a toy example: $\mathcal{B}$ is the decision boundary of binary classification problem, example $\boldsymbol{x}$ lies in a 2-dimensional subspace $\mathcal{S}$ of a 3-dimensional space $\mathbb{R}^{3}$. $\delta$ and $\delta_{\mathcal{S}}$ points in the distance to the nearest decision boundary in 3D space and 2D space respectively. Due to the dimensionality constraint, the attack in lower dimensionality requires a higher perturbation norm budget, as a result, a lower (improvement in) ASR. This intuition could be verified through our analytical tool, Decision Surface, as depicted in Fig. 10(b) where black arrows indicate the distance to cross the decision boundary along the perturbation direction. Notably, inputs in lower dimensionalities require longer distances to traverse the boundary.
IV-E Effect of Model Architecture
In this section, we are interested in the relationship between attack performances and model architectures. It is well-known that different classification models employ distinct architectural designs as well as different decision boundaries. In query-based attacks, we focus on the vulnerability to input perturbations of different architectures. We specifically compare the difficulty of crafting successful adversarial samples among various target model architectures from different families4. In transfer-based attacks, our experiment intends to explore how and what kind of misalignment between surrogate models and target models undermines transferability. We explore transferability between models within and between different families [^4] A family contains variants of the same architecture, e.g., ResNet-34/50/152 in ResNet family, Swin-Tiny/Base/Large in Swin family. , i.e., intra/cross-family transferability.
Figure 11: Vulnerability of different model architectures. The ratio of AQN to ASR of decision-based attacks (left) and score-based attacks (right) versus various target model architectures. An $\infty$ ratio indicates an ASR of 0, $i.e.$, no image is successfully perturbed, hence implying an extremely hard attacked model.
IV-E1 Vulnerability of different model architectures
We choose to implement query-based attacks on model architectures from different families: ResNet-50 from ResNet family [43] which introduces the identity shortcut connection, ${}^{\dagger}$ from VGG family [44] which is characterized by its multiple layers of small convolutional filters with consistent receptive fields, Inception-V3 [45] which uses "inception modules" to effectively grasp both local and broad image details, ConvNeXt-Tiny (ConvNeXt-T) from ConvNeXt family [57] which merges depth-wise and point-wise convolutions using grouped convolutions to boost its ability to understand local and wide-ranging dependencies, ViT/B-16 from ViT family [55] which adopts self-attention and a Transformer framework to perceive global contexts and manage distant dependencies. We utilize the ratio of AQN to MQN as a metric to measure the difficulty. The higher the ratio is, the harder the model is to be attacked. For both decision-based black-box attacks and score-based black-box attacks, we employ untargeted attacks using $\ell_{\infty}$ norm and the ImageNet dataset.
From Fig. 11, we observe that, for the majority of attack methods, attacking SOTA models (ConvNeXt-T and ViT/B-16) tends to yield higher ratio values ($i.e.$, more queries but lower ASR), indicating they are preferable in terms of robustness. Notably, Inception-V3 also demonstrates a comparable or occasionally higher level of resilience against attacks like Sign-OPT, Bandits, AdvFlow and Parsimonious attacks. Conversely, ResNet-50 and ${}^{\dagger}$ prove more susceptible to adversarial perturbations, reflected by their lower ratio values.
To investigate the possible hidden reasons, we use our analysis tool, dbViz, to visualize the decision boundaries of various target models. From the visualization results in Fig. 12, we can see that for models with relatively easy architectures like ResNet-50 and ${}^{\dagger}$, the decision boundaries on the ImageNet dataset are quite complex and fragmented. Looking at the three samples we took for the graphing, some of these samples are located on the edge of the decision boundaries of many categories or are surrounded by the decision boundaries of several categories. In comparison, the decision boundaries of the more complex structures, Inception-V3, ConvNeXt-T and ViT-B/16, appear simpler. In other words, the decision boundaries of more complex models are relatively more intact and continuous, without fragmentation. This is bad news for adversarial attacks, as it means that often a larger attack budget is needed to push a clean sample across the decision boundary of the ground-truth category to become an adversarial sample that meets the constraints. In conclusion, our observation and analysis of decision boundary visualizations across various models suggest that the completeness and continuity of decision boundaries in target models constitute a significant factor contributing to the challenge of successful attacks.
Figure 13: Intra/Cross-family transferability. Matrix of ASR in RAP method[29] (upper left), matrix of DI2-FGSM method[25] (upper middle), and matrix of Adversarial Divergence (upper right), where each entry corresponds to a surrogate-target pair. The y-axis represents surrogate models and the x-axis represents target models. We choose three variants in every model family for our evaluations. Matrices in the second row summarize the first row by averaging entries w.r.t. family. Results of more methods which display a similar phenomenon could be found in Appendix G-A.
IV-E2 Intra/Cross-family transferability
We investigate intra/cross-family transferability via four families using two mechanisms [^5] Models with the same mechanism means they share a common foundational block, e.g., both Vit and Swin benefit from self-attention, as well as ResNet and ConvNeXt from convolution.: two CNNs (classic: ResNet[43], SOTA: ConvNeXt[57]) and two SOTA Transformers (ViT[55] and Swin[56]). ResNet family are rather simple convolutional models with low capacities. We consider ResNet-34/50/152. ConvNeXt family are fully convolutional architectures with designs popularised by Transformers. We consider ConvNeXt-Tiny/Base/Large. ViT family modify the original Transformer to make it work with images and induce inductive biases by some training recipes and design choices. We consider ViT-B/16, ViT-B/32 and ViT-L/16. Swin families are hierarchical Transformers employing a shifting window scheme. We consider Swin-Tiny/Base/Large. We appeal to 12 model architectures in total, leading to 12 × 12 = 144 surrogate-target pairs. Fig. 13 reports intra/cross-family transferability rates for various pairs under the untargeted, $\ell_{\infty}$ attack on ImageNet.
Rates $(i,j),|i-j|\leq 2$, highlighted by yellow bounding boxes, indicate the proportion of adversarial samples crafted on the white-box model $i$ misclassified by the black-box model $j$ within the same family, representing intra-family transferability. Firstly, the results on the diagonal where surrogates precisely align with targets show that SOTA Transformers and CNNs do not provide additional security than classic CNNs. None of them could resist white-box attacks. Therefore, how well an attack would perform is heavily dependent on the surrogate-target alignment rather than the target model architecture. Surprisingly, we observe that all models are vulnerable to intra-family transferability in a non-negligible manner, indicating that surrogates and targets shift with respect to each other but within the same family wouldn’t dramatically undermine transferability.
Rates $(i,j),|i-j|>2$ indicate the proportion of adversarial examples generated to mislead white-box model $i$ that also mislead black-box model $j$ ($i,j$ are from different families), i.e., cross-family transferability. The phenomenon among cross-family transferability is more heterogeneous than intra-family transferability. For the reason behind the heterogeneity, as CNNs induce convolutional inductive biases while Transformers communicate globally via self-attention modules, one intuitive conjecture is that the misalignment that would destroy transferability comes from performing different mechanisms, as previously indicated by [63]. However, the second row in Fig. 13 argues that it’s not an absolute rule. For example, the most transferable family to ResNet is intra-mechanism ConvNeXt, but a high cross-mechanism transferability occurs between ConvNeXt and Swin. Thus, the employed mechanism is not as significant a factor in cross-family transferability as we conjectured.
Instead, we hypothesize that the transferability is implied in the resemblance between adversarial examples generated by the surrogate model and target model, which could be measured by the Adversarial Divergence metric (see Sec. III-C4). Using the analysis tool CAD, we train the conditional adversarial distributions of these models and calculate the Adversarial Divergence of each pair. The results shown in Fig. 13(c) align closely with the transfer success rate illustrated in Fig. 13(a) and (b), $i.e.$, the pair with higher ASR will report a lower Adversarial Divergence value. This alignment verifies our hypothesis that two models with high transferability share a vulnerability to a similar pattern of adversarial perturbation. Additionally, with the effective measurement of misalignment between the surrogates and targets, Adversarial Divergence could guide the search for successful adversarial perturbations and shed light on novel transfer-based attack strategies.
Figure 14: Perturbation norm budget. From left to right: attack results of score-based methods, transfer-based methods from data and optimization perspectives, and transfer-based methods from feature and model perspectives versus the perturbation norm budget $\epsilon$.
IV-F Effect of Attack Budget
In this section, we take a closer look at the impact of the attack budgets. We have observed a decrease in the effectiveness or even complete failure of some methods when they attack on a budget that was not fully considered in their original papers. In the following, we provide explanations for these findings. More specifically, we explore attacks with different perturbation norm budgets and query number budgets.
IV-F1 Perturbation norm budget
In transfer-based attacks, Fig. 14 (middle and right) shows a clear rise in ASR as the perturbation norm budget $\epsilon$ increases across all methods. Interestingly, we find the slope in ASR versus budgets remains fairly consistent across most methods, as evidenced by these parallel lines, indicating these attacks are equally sensitive to the budget. However, exceptions to this consistent pattern are observed in FIA[33], NAA[34], DRA[37] and IAA[38], as their performance drops more significantly in cases of low budgets. Recall that the motivations behind these methods, FIA and NAA both decrease instinct features and increase unimportant features, DRA and IAA both drag images away from the ground-truth data distribution. The common nature implied in these four attacks is to craft meaningful noises that make the images do not "look like" the ground-truth class. Unlike other attacks that greedily perturb the images to increase the classification loss and generate irregular noises, this nature necessitates a higher norm budget to maintain their superiority. This phenomenon could also be observed from model attention perspective via our analytical tool, FullGrad, as visualized in Fig. 15. For instance, consider an image from the "dam" category. With a low norm budget, LGV easily, greedily misleads the target model into a really closed but different category "groyne". DRA, IAA, FIA, NAA fail to alter the features with this small budget, the target model still focuses on the dam’s features and recognizes it correctly. However, when a higher norm budget is applied, these four methods do promote other trivial features. Accordingly, the target model shifts its focus away from the dam and directs its attention to other features. We use IAA as an illustrative example to substantiate the hypothesis that the ability to generate semantic adversarial perturbations requires a relatively high norm budget in Fig. 16. As the budget increases, the perturbation gradually contains more and more obvious features of the "bubble" category. Then this meaningful perturbation overshadows the original features and dominates the classification, as visualized in the attention maps. On the other hand, while this characteristic constrains the attack power in low $\epsilon$ cases, it could make attacks stand out when a high budget is available.
In query-based attacks, we only present the experimental results for score-based attacks. This is because the attack process for score-based attacks starts with a clean sample and introduces adversarial perturbations to create an adversarial sample that crosses the decision boundary of the ground-truth category within a specified $\ell_{p}$ norm perturbation ball. We believe that the choice of different sizes of perturbation balls will also affect the attack performance of different algorithms, which can be considered as one of the attack budgets for score-based attacks. For decision-based attacks, since the attack mechanism starts with an adversarial sample and gradually narrows the distance to the clean sample, the perturbation ball similar to the attack budget in score-based attacks is no longer applicable. For each method in Fig. 14(left), ASR increases with the perturbation budget. For instance, Parsimonious Attack[8] starts with an ASR of 0.27 at the smallest perturbation budget of 1/255 and reaches an ASR of 1.00 at a budget of 32/255. Similarly, the ASR for Square Attack[9] increases from 0.31 to 1.00 as the perturbation budget grows. PPBA[10], NES[11], and Bandits[12] also show a trend of incrementally higher ASR with larger perturbation budgets. At the lowest budget of 1/255, the highest ASR achieved by any method is 0.31 for MCG[19], and at the highest budget of 32/255, the lowest ASR is 0.79 for NES[11], while all other methods reach ASR of 1.00, indicating that expanding the attack budget indeed improves the performance of the attack methods, aligning with our intuitive judgment. It can be seen that $\mathcal{CG}$-attack[18] and MCG[19] perform better than other methods across any attack budget, which once again proves that leveraging additional information from a surrogate model can bring significant improvements to the performance of score-based attacks. Even with a lower attack budget, these two methods maintain a high ASR. Furthermore, methods like Square Attack[9], AdvFlow[15], SignHunter[14], and PPBA[10] also maintain good performance, with their unique attack mechanisms (for example, Square Attack[9] explores the perturbation among its $L_{\infty}$ vertices, and AdvFlow[15] introduces a flow model as the perturbation sampling model, etc.) providing references for the proposal of new methods.
Figure 19: Momentum-based attack results. (a) White-box(dashed) and black-box(solid) ASR w.r.t. iteration numbers. (b) Black-box ASR w.r.t. step size. Lines w/ markers plot momentum-based methods ASR versus step size with 50 iterations(more than the recommended 10 iterations), where the solid marker indicates ASR with unadapted step size $\alpha^{*}$. Lines w/o markers show ASR with 10 iterations and step size $\alpha^{*}$.
IV-F2 Query number budget
In query-based attacks, our research focuses on both decision-based black-box attacks and score-based black-box attacks.
For decision-based black-box attacks, we will show the evaluation of Mean Squared Error (MSE) between the optimized adversarial examples and the clean example. We will vary the number of queries while ensuring a 100% attack success rate. This metric represents the magnitude of the adversarial perturbation. A smaller value indicates that the adversarial example is gradually approaching the clean example to satisfy the adversarial constraint. From the results in Fig. 19, we can see the MSE of the PSBA attack drops most significantly when the query number is still small. Additionally, NonLinear-BA and QEBA, which inherit the HSJA attack framework, also show promising attack performance. These two methods exhibit similar attack behavior, with comparable rates of perturbation magnitude reduction as the number of queries increases. On the other hand, the attack performance of Sign-OPT and OPT is less satisfactory, as the MSE remains high even after a significant number of queries. Furthermore, SFA and Rays rapidly reduce the MSE to lower values at a relatively low query cost.
For score-based black-box attacks, we will show how the ASR changes with an increasing number of queries. We evaluate all decision-based black-box attacks and score-based black-box attacks on ResNet-50 using $\ell_{\infty}$ norm untargeted attacks, with the ImageNet dataset as the test set. For score-based black-box attacks, we observe from Fig. 19 that MCG and $\mathcal{CG}$-attack significantly outperform other methods, achieving the highest ASR levels at the fastest rates. Among the gradient-estimation-based methods, SignHunter is noteworthy as it quickly boosts the ASR to the highest level with minimal query cost. In fact, under untargeted attacks, we can see from the figure that almost all methods achieve a high ASR. It is worth noting that AdvFlow, although it eventually achieves a high ASR, incurs a significant cost as evidenced by its nearly vertical line, indicating a high number of queries for each clean example. Meanwhile, among the random search methods, the Square attack performs exceptionally well, achieving the highest ASR at the fastest rate.
In transfer-based attacks, the term query number budget refers to the number of iterations of the white-box attack against the surrogate model. Due to the iterative nature of transfer-based attacks in crafting adversarial examples, it’s therefore necessary to investigate the influence of the number of iterations on the efficacy of these methods. Fig. 17 visualizes the results of untargeted, $\ell_{\infty}$ attack on ImageNet, where all method shares a common maximum perturbation $\epsilon^{*}$ and step size $\alpha^{*}$. We see that basically optimal performance is attained with 50 iterations, and plateaus with further iterations for data-perspective, feature-perspective, model-perspective, and part of optimization-perspective methods. However, it is worth noting that this stability does not extend to MI [26], NI [23], PI [27], and their corresponding composite attacks (VMI and VNI). They perform best with exactly 10 iterations as suggested in their original papers, however, deteriorate for additional rounds. Recall that the fundamental concept shared by MI, NI and PI is momentum-based gradient optimization. However, inadequate step size will impede the progress of adversarial examples towards the optimal solution. That is to say, over multiple iterations, accumulated momentum may lead to oscillating around the optimal solution or getting stuck in local minima, resulting in a reduction in white-box success rate and, consequently, a decrease in black-box ASR, as shown in Fig. 19(a). Based on this observation, in Fig. 19(b), we propose to adapt the step size of momentum-based attacks. It is evident that even with a suboptimal 50 iterations, a gradual adaptation of step size enables ASR to approach or surpass the optimal results achievable with 10 iterations. These findings boil down to that momentum-based methods’ transferability heavily relies on the early stop, and essentially, appropriate step sizes.
IV-G Effect of Defense
@@
Figure 20: Different query-based attacks against defenses. Each tick on the horizontal axis represents the performance of each method of decision-based attack (top) and score-based attack (bottom) under different defenses. We use (shaded) bars to represent the MQNs (AQNs) and values above bars to represent ASRs.
@@
Figure 21: Intra/Cross-training-strategy transferability. The matrices in the first column represent the ASRs for DI2-FGSM (top) and RAP (bottom) attacks, with each entry corresponding to a surrogate-target pair. The y-axis and x-axis represent surrogate models and target models, respectively. We employ the regular and AT versions of three model architectures for our evaluations. Additional results of more attacks exhibiting similar trends can be found in Appendix G-B. The other columns show the magnitude of the FSM for six models involved in this analysis. Low frequencies are located at the center, and high frequencies at the corners.
@@
In our previous experimental analyses, we focused on clean models. Now, we shift our focus to attacks on victims involved of defense mechanisms. Here we consider defenses mentioned in Sec. III-B. In the context of query-based attacks, we aim to investigate the evasion capabilities of different attack methods against training-time-based defense and inference-time-based defense and explore the advantages and disadvantages of these two types of defenses. As for transfer-based attacks, in Sec. IV-E we have observed the strong intra-family transferability and heterogeneous cross-family transferability among clean models, we further expand on the previous analyses and now examine the transferability between normally trained models and adversarially trained models, i.e., intra/cross-training-strategy transferability.
IV-G1 Different query-based attacks against defenses
@@
In this study, we aim to investigate the evasion capabilities of different attack methods against defenses. For our evaluation, we have chosen the ImageNet dataset, which includes two AT models following the ResNet architecture. We specifically evaluate decision-based black-box attacks and score-based black-box attacks using targeted attacks with $\ell_{\infty}$ norm. We will present the AQN, MQN and ASR achieved by different attack methods against the defenses in Fig. 20. It is important to note that some defense methods may increase the query numbers of attack methods, but the corresponding increase in ASR indicates that the defense primarily raises the cost of the attack. In decision-based black-box attacks, methods like OPT and Sign-OPT face challenges under both training-time and inference-time defenses, often failing to attack even one clean example. However, attacks like QEBA, NonLinear-BA, and PSBA can bypass defenses to some extent, achieving ASRs of around 40%. Remarkably, random search methods such as SFA and Rays consistently outperform, with ASRs exceeding 70% across all defense scenarios, indicating that attacks effective on clean models also perform well against defended models. In score-based black-box attacks, there’s a notable variation in attack performance across different methods, with query numbers ranging from a few thousand to 30,000-40,000. Among all defenses, the combination-based method MCG exhibits the highest ASR and the fewest queries, demonstrating only a 10%-15% reduction in ASR compared to attacking ResNet-50.
IV-G2 Training-time-based defense versus inference-time-based defense
@@
Defense methods against adversarial attacks fall into two categories: training-time-based and inference-time-based defenses. Assessing their respective pros and cons is essential. Referring to Fig. 20 (top), we compare AT and RND, as AAA exclusively defends against score-based attacks. The results reveal that RND primarily increases the query cost for attackers to achieve performance comparable to, or slightly lower than, those without defense strategies, leading to a trade-off between performance and query cost. Conversely, the query numbers for attacking AT models are similar to those on the clean model (ResNet-50), while the achieved ASR significantly diminishes, indicating a different kind of resistance. In summary, AT could effectively reduce ASR but at the cost of considerable model training overhead. RND provides lightweight and real-time defense while not significantly reducing ASR.
@@
IV-G3 Intra/Cross-training-strategy transferability

### v2

@@ C-B2 Input dimensionality
We also investigate how the dimensionality the adversary restricted to affects its performance? We observe that adversarial attacks on images of lower dimensions are less successful, and suggest that the constraints of the dimensionality results in a longer distance needed to cross the decision boundary, thus requires larger perturbations to be effective. We now quantify the sensitivity to input dimensionality via conducting untargeted, $\ell_{\infty}$ attack on ImageNet[80] dataset with varying image size. It’s important to note that when working with a clean dataset with downsampled images, the clean accuracy naturally decreases, potentially confusing our analysis. Therefore, we use the improvement in ASR after adding adversarial perturbations on benign examples resized to different dimensionalities. This metric only serves as the measure of the transfer-based attack performance. Query-based attacks define the percentage of finally adversarial images in initially non-adversarial images as ASR and thus are not influenced by the dropped clean accuracy. Fig. 22 shows that all adversaries suffer from a considerable failure rate when facing the robustness existing in low dimensional classification tasks.
This intuition behind this phenomenon is clearly illustrated in Fig. 23 (left). Consider a toy example: $\mathcal{B}$ is the decision boundary of binary classification problem, example $\boldsymbol{x}$ lies in a 2-dimensional subspace $\mathcal{S}$ of a 3-dimensional space $\mathbb{R}^{3}$. $\delta$ and $\delta_{\mathcal{S}}$ points in the distance to the nearest decision boundary in 3D space and 2D space respectively. Due to the dimensionality constraint, the attack in lower dimensionality requires a higher perturbation norm budget, as a result, a lower (improvement in) ASR. This intuition could be verified through our analytical tool, Decision Surface, as depicted in Fig. 23 (right) where black arrows indicate the distance to cross the decision boundary along the perturbation direction. Notably, inputs in lower dimensionalities require longer distances to traverse the boundary.
C-C Effect of Defense
@@
Figure 24: Different query-based attacks against defenses. The bar chart and line chart represent the query number and ASR, respectively, of decision-based attacks (top) and score-based attacks (bottom) against different defenses. We use solid bars to represent the MQNs and shaded bars to represents AQNs. For the two chosen AT models, one is from [97], the other is trained using fast AT.
@@
Figure 25: Intra/Cross-training-strategy transferability. The matrices in the first column represent the ASRs for DI2-FGSM (top) and RAP (bottom) attacks, with each entry corresponding to a surrogate-target pair. The y-axis and x-axis represent surrogate models and target models, respectively. We employ the regular and AT versions of three model architectures for our evaluations. Additional results of more attacks exhibiting similar trends can be found In Fig. 26. The other columns show the magnitude of the FSM for six models involved in this analysis. Low frequencies are located at the center, and high frequencies at the corners.
@@
In our previous experimental analyses, we focused on clean models. Now, we shift our focus to attacks on target models involved of defense mechanisms. Here we consider defenses mentioned in Supplementary A-4. In the context of query-based attacks, we aim to investigate the evasion capabilities of different attack methods against training-time defense and inference-time defense. We also conduct an in-depth discussion on the advantages and disadvantages of these two types of defenses. In decision-based black-box attacks, random search methods like SFA [48] and Rays[49] perform the best, achieving ASRs over 70%, while traditional methods such as OPT [50] and Sign-OPT [51] struggle under defenses. In score-based attacks, combination-based methods like MCG [42] stand out with the highest ASR and fewest queries, maintaining strong performance even when defenses are applied. Besides, we can also observe that training-time defense methods (e.g., adversarially trained models) effectively suppress the attack success rates of various methods, while inference-time defense methods primarily focus on increasing the number of queries required for each attack. As for transfer-based attacks, in Sec. IV-D2, we have observed the strong intra-family transferability and heterogeneous cross-family transferability among clean models, we further expand on the previous analyses and now examine the transferability between normally trained models and adversarially trained models, i.e., intra/cross-training-strategy transferability. It is observed that stronger transferability occurs within models that share the same training strategies, rather than between different strategies. This is attributed to the different spatial frequency features used by the models for predictions.
C-C1 Different query-based attacks against defenses
@@
In this study, we aim to investigate the evasion capabilities of different attack methods against defenses. For our evaluation, we have chosen the ImageNet dataset, which includes two AT models following the ResNet architecture. We specifically evaluate decision-based black-box attacks and score-based black-box attacks using targeted attacks with $\ell_{\infty}$ norm. We will present the AQN, MQN and ASR achieved by different attack methods against the defenses in Fig. 24. It is important to note that some defense methods may increase the query numbers of attack methods, but the corresponding increase in ASR indicates that the defense primarily raises the cost of the attack. In decision-based black-box attacks, methods like OPT [50] and Sign-OPT [51] face challenges under both training-time and inference-time defenses, often failing to attack even one clean example. However, attacks like QEBA [54], NonLinear-BA [55], and PSBA [56] can bypass defenses to some extent, achieving ASRs of around 40%. Remarkably, random search methods such as SFA [48] and Rays [49] consistently outperform, with ASRs exceeding 70% across all defense scenarios, indicating that attacks effective on clean models also perform well against defended models. In score-based black-box attacks, there’s a notable variation in attack performance across different methods, with query numbers ranging from a few thousand to 30,000-40,000. Among all defenses, the combination-based method MCG [42] exhibits the highest ASR and the fewest queries, demonstrating only a 10%-15% reduction in ASR compared to attacking ResNet-50.
C-C2 Training-time defense vs. inference-time defense
@@
Defense methods against adversarial attacks fall into two categories: training-time and inference-time defenses. Assessing their respective pros and cons is essential. Referring to Fig. 24(top), we compare AT and RND [99], as AAA [100] exclusively defends against score-based attacks. The results reveal that RND primarily increases the query cost for attackers to achieve performance comparable to, or slightly lower than, those without defense strategies, leading to a trade-off between performance and query cost. Conversely, the query numbers for attacking AT models are similar to those on the clean model (ResNet-50), while the achieved ASR significantly diminishes, indicating a different kind of resistance. In summary, AT could effectively reduce ASR but at the cost of considerable model training overhead. RND provides lightweight and real-time defense while not significantly reducing ASR.
@@
C-C3 Intra/Cross-training-strategy transferability